OpenStudy (anonymous):

If f(x)= (x^2)/2 (lnx)-(3/4) x^2 + C and f(1)=6, then find the values of C.

OpenStudy (freckles):

replace x with 1

OpenStudy (freckles):

then replace f(1) with 6

OpenStudy (freckles):

\[f(1)=\frac{1^2}{2} \ln(1)-\frac{3}{4}(1)^2+C \\ 6=\frac{1^2}{2} \ln(1)-\frac{3}{4} (1)^2+C\]

OpenStudy (freckles):

simplify a bit and then solve for C

OpenStudy (anonymous):

I came down to (1/2) ln - (3/4) + C

OpenStudy (freckles):

ln(1) is 0

OpenStudy (freckles):

do you know how to solve 6=-3/4+C for C?

OpenStudy (anonymous):

Oh. So would C=27/4?

OpenStudy (freckles):

yes

OpenStudy (anonymous):

Oh. Okay. Thank You!

OpenStudy (freckles):

np